OpenStudy (idealist10):

How to integrate dz/(z^2+2z/x)?

OpenStudy (idealist10):

@ganeshie8 @thomaster @phi @e.mccormick @kropot72

OpenStudy (idealist10):

@aum

OpenStudy (idealist10):

@wio @zepdrix

OpenStudy (idealist10):

A/z+B/(z+2/x)=1 A(z+2/x)+Bz=1 A=x/2, B=-x/2

zepdrix (zepdrix):

\[\Large\rm \int\limits \frac{dz}{z^2+2\frac{z}{x}}\]

zepdrix (zepdrix):

Oh you decided to go with Partial Fractions? Fair enough. :p

OpenStudy (idealist10):

Sorry, I can't see what you wrote since processing math is 0%. Can you please take a snapshot?